A garden hose is originally full of motionless water. What additional force is necessary to hold the nozzle stationary after the water flow is turned on if the discharge rate is 0.600 kg/s with a speed of 25.0 m/s?
1
Expert's answer
2020-09-30T12:29:23-0400

We can find the additional force from the formula:


"F = v \dfrac{dm}{dt},"

here, "v = 25 \ \dfrac{m}{s}" is the speed of water, "\dfrac{dm}{dt} = 0.6 \ \dfrac{kg}{s}" is the discharge rate of water.

Then, we get:


"F= 25 \ \dfrac{m}{s} \cdot 0.6 \ \dfrac{kg}{s} = 15 \ N."

Answer:

"F = 15 \ N."